OpenStudy (skatingdj):

Hi, can somebody help me make some equations? I need 3, one true, one false, and one open that includes 2-3 steps. I know what it means but I don't know how to make a decent one, it's supposed to be for my classmates to review and I'm not good at making them up:p Any help is appreciated.

563blackghost (563blackghost):

It depends on what you are learning for the Unit. `Equation is a statement with values that include an equal sign` could be anything. Please specify :)

OpenStudy (skatingdj):

Sorry about that, it's simple Algebra. Using PEMDAS, short equations. Does that help?

jhonyy9 (jhonyy9):

same with this 3x = (6x)/2 3x = (2*3x)/2 3x = 3x so tis is true second x= 8x/4 x = (2*4x)/4 x = 2x this is false third 3x +4 = 0 subtract from both sides 4 and will get 3x = -4 divide both sides by 3 and will get x = -4/3 hope helped

jhonyy9 (jhonyy9):

@OtherWorldly sorry but i not can reply to you just in this style bc. my submitt button is blocked from when my past champion title profile was deleted yes i know something spanish - why ?

OpenStudy (otherworldly):

yeah i totally forgot

jhonyy9 (jhonyy9):

Skating not id right what i ve wrote above ?

jhonyy9 (jhonyy9):

not is sorry

jhonyy9 (jhonyy9):

or where you dont understand it ?

OpenStudy (skatingdj):

No, I knew what the question was asking, but I'm just really not good at creating equations! XD But thank you so much for the help!!!